OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Show how to solve this using any method.

OpenStudy (anonymous):

x=98 y=72

OpenStudy (anonymous):

x+y=170 x-y=26 therefore x=170-y subs in eq 2 170-y-y=26 -2y=-144 y=72 x=98